Homework 1 11-721: Grammars and Lexicons Fall term, 2003

Due Monday, September 8

1 Diagnostic Tests for Parts of Speech

1.1 Educational Goals of the Exercise (What you are supposed to learn)

• In this exercise you will continue our in-class discussion on using tests to distinguish parts of speech. The exercise focuses on distinguishing adjectives from prepositions.

• The goal of the exercise is for you to learn to apply tests and interpret the results.

• A secondary goal of the exericise is for you to think about linguistic methodology. Is it always possible to formulate reliable tests? Are parts of speech really well-defined? Or are they ”squishy” categories?

1.2 Formulation of the Tests

1. Comparative Test: Adjectives can be made into comparatives. Monosyllabic adjectives typically take -er and -est. Multi-syllabic adjectives take more and most.

Non-controverisal examples: Adjectives in comparative forms:

(1) a. He is taller.

b. You have to be 48 inches tall to ride this roller coaster. Shorter kids can’t ride.

c. He is more interesting.

d. We should avoid more problematic theories.

Other parts of speech (nouns, verbs, prepositions) don’t form comparatives.

2. Very Test: Adjectives (and adverbs) can be modified by very. Other parts of speech cannot be modified by very.

(2) a. He is very tall. b. The book is very interesting. c. I read the very interesting book.

3. Right Test: Prepositions can be modified by the intensifier right.

(3) a. The book is right on the table. b. Look right behind the couch. c. The school is right across the street.

Now consider the following two sentences:

(4) a. This is worth a dollar. b. Sam is near the fence.

• Task 1: Apply the comparative test to worth and near.

• Task 2: Apply the very test to worth and near.

• Task 3: Apply the right test to worth and near.

• Task 4: Can you conclude whether worth is an adjective or preposition? Why or why not? (Results may vary according to your grammaticality judgements and creativity in applying and altering tests.)

• Task 5: Can you conclude whether near is an adjective or preposition? Why or why not? (Results may vary according to your grammaticality judgements and creativity in applying and altering tests.)